OpenStudy (anonymous):

Randall bought a bond with a face value of $6000 and a coupon rate of 7.25%. The bond will mature in 5 years. How much interest will he receive semiannually

OpenStudy (wolf1728):

Total = Principal * (1+rate)^years Total = 6,000 * 1.0725^5 Total = 6,000 * 1.4190134251 Total = 8,514.08 For compounded interest, the formula is If compounded semi-annualy the rate is: Total = Principal * (1+rate/n)^n*years where n is the number of compounding periods which for semi annual means two. Total = 6,000 * (1+.0725/2)^2*5 Total = 6,000 * (1.03625)^10 Total = 6,000 * 1.4277278645 Total = 8,566.37 Perhaps you meant the interest received every 6 months 5 year interest = 2,514.08 every 6 months = 2,514.08 = 251.41